### v2

# Koopman Autoencoders with Continuous-Time Latent Dynamics for Fluid Dynamics Forecasting

Abstract
Learning surrogate models for time-dependent PDEs requires balancing expressivity, stability, and computational efficiency. While highly expressive generative models achieve strong short-term accuracy, they rely on autoregressive sampling procedures that are computationally expensive and prone to error accumulation over long horizons. We propose a continuous-time Koopman autoencoder in which latent dynamics are governed by a parameter-conditioned linear generator. This formulation enables exact latent evolution via matrix exponentiation, allowing predictions at arbitrary temporal resolutions without autoregressive rollouts. We evaluate our method on challenging fluid dynamics benchmarks and compare against autoregressive neural operators and diffusion-based models. We evaluate our method on challenging fluid dynamics benchmarks against autoregressive neural operators and diffusion-based models. Our results demonstrate that imposing a continuous-time linear structure in the latent space yields a highly favorable trade-off: it achieves massive computational efficiency and extreme long-horizon stability while remaining competitive in short-term generative accuracy.
1 Introduction
Simulating turbulent flows is a fundamental problem in fluid dynamics, with applications ranging from aerodynamics and weather prediction to engineering design. Traditional computational models such as Direct Numerical Simulation (DNS), Large-Eddy Simulation (LES), and Reynolds-averaged Navier–Stokes (RANS) (Ghasemian et al., 2017; Slotnick et al., 2014) are extremely expensive in both computation and memory, especially for long time horizons or high-resolution domains. This motivates the development of data-driven surrogate models that can efficiently forecast complex unsteady flows.
Deep learning has recently emerged as a promising tool for flow prediction (Borrelli et al., 2022; Srinivasan et al., 2019; Fukami et al., 2020; Morimoto et al., 2021). Classical data-driven solvers operate directly on the data space, mapping flow fields at the current timestep to the next autoregressively using models such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012; Bhatnagar et al., 2019; Guo et al., 2016), Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Rubanova et al., 2021; Pfaff et al., 2020) and Neural Operators (Lu et al., 2021; Li et al., 2021). While such data-space models excel at capturing fine-scale features, they tend to suffer from error accumulation (Brandstetter et al., 2022), high computational cost due to step-by-step rollouts, and reduced stability over long horizons.
Latent-space models compress high-dimensional flow fields into low-dimensional representations and evolve them over time using learned dynamics. Prior work (Vlachas et al., 2018; Wu et al., 2022) demonstrates that learning global latent dynamics can provide a simple and scalable alternative. A principled formulation of latent dynamics is given by Koopman theory (Koopman, 1931), which represents nonlinear systems through linear evolution in a lifted feature space. In practice, this operator is infinite-dimensional and must be approximated. Koopman autoencoders address this by learning a nonlinear mapping to a finite-dimensional latent space in which the dynamics evolve linearly (Takeishi et al., 2017; Lusch et al., 2018).
However, existing Koopman-based approaches typically learn discrete-time operators that advance the latent state by a fixed timestep, tying the learned dynamics to the temporal resolution of the training data. Inspired by continuous-time modeling approaches such as neural ODEs (Chen et al., 2018), we instead represent latent evolution using a linear Koopman generator. Unlike general neural ODE models, this formulation preserves a linear latent structure, enabling closed-form solutions through matrix exponentiation, interpretable spectral properties of the dynamics, and efficient long-horizon forecasting without iterative integration. In addition, we model this continuous operator as a function of governing physical parameters (e.g., Reynolds and Mach numbers), which enables a single model to capture families of dynamical systems across different regimes rather than learning a fixed operator tied to a single configuration.
Our approach focuses on deterministic latent dynamics designed to prioritize long-horizon stability, temporal flexibility, and computational efficiency. To rigorously evaluate these capabilities, we benchmark our model against a diverse set of baselines, including expressive generative models and data-space autoregressive predictors. Our evaluation demonstrates that enforcing a structured, continuous-time linear operator in the latent space yields significant advantages over highly expressive but unstable autoregressive methods. Ultimately, these findings establish continuous-time Koopman models as a highly scalable and robust alternative for long-horizon PDE forecasting.
Contributions:
  • Continuous-Time Koopman Autoencoder: We formulate latent evolution as a continuous-time linear dynamical system governed by a learned Koopman operator, enabling closed-form integration via matrix exponential and eliminating the need for autoregressive rollouts.
  • Physics-Informed, Parameter-Conditioned Latent Dynamics: The Koopman operator is conditioned on physical parameters, allowing a single model to represent families of flow regimes while maintaining physically consistent predictions.
  • Temporal Resolution Super-Resolution: By modeling the latent dynamics in continuous time, the learned system can be evaluated at arbitrary timesteps, enabling zero-shot prediction at unseen temporal resolutions.
  • Efficient Long-Horizon Forecasting: The proposed formulation enables non-autoregressive prediction through matrix exponential evolution of the latent state, resulting in substantially faster inference compared to diffusion-based baselines.
2 Related Work
Autoregressive prediction in data space. Classical deep learning approaches for flow forecasting adopt autoregressive prediction in the data space. CNNs, particularly U-Net architectures (Ronneberger et al., 2015), have been widely used to capture complex spatial features in a dense image modeling framework (Deo et al., 2023; Wang et al., 2019). Other approaches have also explored generative models. For instance, Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) have been employed to predict complex unsteady flow dynamics around objects like cylinders (Lee and You, 2019). Similarly, conditional diffusion models (Batzolis et al., 2021) have been benchmarked for their ability to predict turbulent flow simulations autoregressively (Kohl et al., 2026).
Latent-space autoregressive models. Many prior works compress flow fields into latent representations and evolve them using autoregressive dynamics. While this decouples spatial and temporal complexity, these approaches rely on step-to-step prediction, leading to error accumulation and limited temporal flexibility. Classical examples include CNN autoencoders coupled with LSTMs (Hochreiter and Schmidhuber, 1997) and Recurrent Neural Networks (RNNs) to propagate the latent state forward in time (Eivazi et al., 2020; Hasegawa et al., 2020; Bukka et al., 2020; Maulik et al., 2021; Zhou and Cheng, 2025). Similarly, Graph Neural Networks (GNNs) have proven effective for systems with irregular geometries to conduct temporal evolution in latent space (Sanchez-Gonzalez et al., 2020). More recently, transformers (Vaswani et al., 2017) have also been adapted for this task, leveraging its powerful attention mechanism to capture long-range temporal dependencies in the latent dynamics of fluid flows (Hemmasian and Farimani, 2023).
Koopman autoencoders. Building upon Lusch et al. (2018)’s foundational work, subsequent studies have introduced various enhancements. Azencot et al. (2020) proposed Consistent Koopman Autoencoders, integrating both forward and backward dynamics to enhance accuracy for long-time predictions. The Temporally-Consistent Koopman Autoencoder (tcKAE) (Nayak et al., 2025) achieved accurate long-term predictions with limited and noisy training data through a consistency regularization term. Halder et al. (2026) proposed Koopman $\beta$-variational autoencoders for reduced-order modelling (ROM) of turbulent flows, while $K^{2}$VAE integrates KalmanNet with KAEs to refine predictions and model uncertainty (Wu et al., 2025). Similar to our work, Menier et al. (2025) modeled continuous-time latent dynamics using a linear Koopman operator augmented by a nonlinear Mori-Zwanzig closure.
Different from existing works, we formulate the latent dynamics as a strictly continuous-time linear ODE. Crucially, we parameterize the Koopman operator as a function of governing physical inputs. This decoupled, continuous formulation allows for highly efficient integration at arbitrary temporal resolutions via matrix exponentiation, while the physical conditioning enables a single model to adapt and generalize seamlessly across different flow regimes.
3 Background
3.1 Koopman Autoencoders
Koopman operator theory provides a linear perspective on nonlinear dynamics. Consider a discrete-time dynamical system with states $x_{t}\in\mathbb{R}^{N_{d}}$, where $x_{t}$ denotes the state of the system at time $t\in N^{d}$. The state evolves to its next timestep $t+\Delta t$ according to a flow map $f:\mathbb{R}^{N_{d}}\to\mathbb{R}^{N_{d}}$:
$x_{t+\Delta t}=f(x_{t}).$ (1)
Koopman operator theory shifts the focus from the state variables $x_{t}$ to a set of measurement functions $h:\mathbb{R}^{N_{d}}\to\mathbb{R}$, which are elements of an infinite-dimensional Hilbert space. The Koopman operator $\mathcal{K}$ acts on measurement functions $h$ linearly with the dynamics:
$(\mathcal{K}h)(x_{t})=h(f(x_{t}))=h(x_{t+\Delta t}).$ (2)
In this formulation, the underlying nonlinear system (1) is transformed into a linear evolution in the infinite-dimensional space of observables. The challenge lies in choosing or constructing a suitable set of observables. KAEs address this by learning the observables from data, training an encoder which maps the state into a latent representation $\mathcal{E}:\mathbb{R}^{N_{d}}\to\mathbb{R}^{N_{z}}$,
$z_{t}=\mathcal{E}(x_{t}),$ (3)
which serves as a finite-dimensional approximation of the observable space. In this latent space, the evolution can be modelled as an approximately linear finite-dimensional Koopman operator $\mathbf{K}$,
$z_{t+\Delta t}\approx\mathbf{K}z_{t}.$ (4)
A decoder $\mathcal{D}$ then reconstructs the state from latent variables,
$\hat{x}_{t+\Delta t}=\mathcal{D}(z_{t+\Delta t}).$ (5)
While most KAE formulations operate in discrete time, directly mapping $z_{t}\to z_{t+\Delta t}$, such updates are tied to the temporal resolution of training data. This dependence can hinder generalization across timesteps and reduce robustness in scenarios with different or irregular temporal resolutions. To overcome this limitation, we propose continuous-time variants which treat latent dynamics as an ODE:
$\frac{dz}{dt}=\mathbf{K}z,$ (6)
enabling evaluation at arbitrary timesteps and extrapolation beyond the training discretization.
4 Method
In this section, we present the methodology of our proposed continuous-time KAE. We begin by defining the network architecture. We then detail the mathematical formulation of our proposed framework. Finally, we outline the training strategy.
4.1 Proposed Model
The proposed KAE consists of three main components: a dual-stream Transformer encoder, a parametric latent Koopman operator, and a CNN-based decoder. More details on the exact implementation can be found in Appendix B.1.
Let $x_{t}\in\mathbb{R}^{N_{d}}$ denote the high-dimensional physical state of the system. In our training datasets, the flow snapshots are provided at discrete timestamps $t_{i}$, uniformly spaced by a fixed time interval $\Delta t$, such that $t_{i+1}=t_{i}+\Delta t$. Although the formulation naturally supports irregularly sampled data, we use uniformly spaced timesteps in our experiments to align with standard benchmark protocols and enable direct comparison with existing baselines.
4.1.1 Encoder and Decoder
We employ a Transformer-based encoding architecture consisting of two symmetric streams: a history encoder ($\mathcal{E}_{\text{history}}$) and a present encoder ($\mathcal{E}_{\text{present}}$). To maintain consistency with benchmark evaluations, we formulate the temporal prediction as an Auto-Regressive process of order 2 (AR-2). Mathematically, we model the conditional distribution of the future state as $p(x_{t_{i+1}}|x_{t_{i}},x_{t_{i-1}})$.
Crucially, both encoders share identical weights and architecture, effectively acting as a single Siamese network that processes the immediate past state $x_{t_{i-1}}$ and the current state $x_{t_{i}}$ in parallel. The initial latent state $z_{t_{i}}$ is then constructed by averaging the feature representations from both streams:
$z_{t_{i}}=\frac{1}{2}\left(\mathcal{E}_{\text{present}}(x_{t_{i}})+\mathcal{E}_{\text{history}}(x_{t_{i-1}})\right).$ (7)
A decoder $\mathcal{D}$ reconstructs the physical state from the latent vector, $\hat{x}_{t_{i}}=\mathcal{D}(z_{t_{i}})$.
4.1.2 Koopman Operator
Next, we discuss the structure of the Koopman operator. While conditioning neural surrogate models on physical parameters is common in PDE forecasting, incorporating this conditioning directly into a Koopman operator has received comparatively little attention. Therefore, a key innovation in our approach is explicitly conditioning the continuous-time Koopman operator on these external control parameters, denoted by $\phi$. In our setting, $\phi$ represents physical parameters such as Reynolds or Mach numbers. Unlike traditional Koopman approaches where a single static operator is learned for the entire domain, our operator adapts dynamically based on the specific physical regime.
Our proposed continuous-time model interprets the latent evolution as an ODE,
$\frac{dz}{dt}=\mathbf{K}_{\text{cont}}(\phi)z,$ (8)
where $\mathbf{K}_{\text{cont}}$ governs the linear evolution in latent space. The state at a future time, $t_{i}+\Delta t$, is obtained by integrating this ODE:
$z_{t_{i}+\Delta t}=z_{t_{i+1}}=z_{t_{i}}+\int_{0}^{\Delta t}\frac{\mathrm{d}z}{\mathrm{d}t}dt,$ (9)
where fourth-order Runge–Kutta (RK4) is used to conduct this integration stably.
Finally, we define the structure of the operator. To capture the dependency on physical parameters $\phi$, we express $\mathbf{K}_{\text{cont}}(\phi)$ as the sum of a static base dynamics matrix and a parametric adjustment:
$\mathbf{K}_{\text{cont}}(\phi)=\mathbf{K}_{0}+\mathcal{N}_{\psi}(\phi).$ (10)
Here, the base dynamics are governed by a learnable matrix $\mathbf{K}_{0}$, which captures the global, invariant behavior shared across all flow regimes. $\mathcal{N}_{\psi}$ is a Low-Rank Adaptation (LoRA (Hu et al., 2022)) module parameterized by $\psi$. We use a low-rank parameterization to control the operator spectrum and maintain numerical stability. This restricts $\mathbf{K}_{\text{cont}}$ from deviating too radically from the stable base dynamics $\mathbf{K}_{0}$, while significantly reducing the computational overhead of parameterizing an entire matrix for every distinct flow regime.
During training, we inject small Gaussian noise into the conditioning parameters $\phi$. This acts as a regularizer that improves robustness to sparse parameter sampling and discretization artifacts. In practice, this encourages the learned Koopman operator to vary smoothly across parameter space, improving interpolation between regimes.
4.2 Training
The training procedure follows a recurrent rollout scheme. Given an input sequence context $\{x_{t_{i-1}},x_{t_{i}}\}$, the model predicts a future trajectory of length $N$, denoted as $\{\hat{x}_{t_{i+1}},\hat{x}_{t_{i+2}},\dots,\hat{x}_{t_{i+N}}\}$. An architecture overview is provided in Figure 1.
The training objective is defined as a weighted combination of several loss components that enforce accurate reconstruction, long-horizon prediction, latent-space consistency, and physics-informed regularization. All losses are computed as mean squared errors (MSE) unless stated otherwise.
Reconstruction loss.
The reconstruction loss enforces consistency between the reconstructed initial state produced by the decoder and the corresponding ground-truth state. Let $\mathbf{x}_{t_{i}}=\{x_{t_{i},q}\}_{q\in\mathcal{Q}}$ denote the flow states at time $t_{i}$, decomposed into its physical components (e.g. velocity components, pressure), where $\mathcal{Q}$ indexes the set of state variables shared by the prediction and ground-truth representations. The reconstruction loss is defined as
$\mathcal{L}_{\mathrm{recon}}=\sum_{q\in\mathcal{Q}}\mathbb{E}_{x\sim p(data)}\left[\left\|\hat{x}_{t_{i},q}-x_{t_{i},q}\right\|_{2}^{2}\right],$ (11)
where $x\sim p(data)$ denotes the empirical data distribution induced by the training dataset.
Rollout prediction loss.
The rollout prediction loss penalizes errors over future time steps produced by iterating the learned latent dynamics and decoding the resulting latent states. This term mitigates error accumulation by training on multi-step rollouts.
$\mathcal{L}_{\mathrm{pred}}=\sum_{q\in\mathcal{Q}}\mathbb{E}_{x\sim p(data)}\left[\sum_{j=1}^{N}w_{i+j}\left\|\hat{x}_{t_{i+j},q}-x_{t_{i+j},q}\right\|_{2}^{2}\right],$ (12)
where $N$ is the rollout horizon and $\{w_{i+j}\}_{j=1}^{N}$ are normalized temporal weights. In this work, the weights follow a cosine schedule that emphasizes early prediction accuracy: $w_{i+j}\propto\frac{1}{2}\left(1+\cos\left(\frac{\pi(i+j-1)}{N-1}\right)\right)$. We use a decaying temporal weighting because autoregressive models are highly susceptible to compounding errors; enforcing stricter accuracy on immediate short-term predictions ensures the model establishes a correct base trajectory.
Latent consistency loss.
To encourage consistency between the predicted latent trajectory and the latent representations obtained by encoding the ground-truth states, we include a latent-space loss,
$\mathcal{L}_{\mathrm{latent}}=\mathbb{E}_{x\sim p(data)}\left[\sum_{j=1}^{N}\left\|z_{t_{i+j}}-\mathcal{E}_{\mathrm{present}}(x_{t_{i+j}})\right\|_{2}^{2}\right],$ (13)
where $z_{t_{i}}$ denotes the latent state obtained by rolling out the latent dynamics, and $\mathcal{E}_{\mathrm{present}}(\cdot)$ is the encoder network.
Physics-informed regularization.
In addition to the standard reconstruction and prediction losses, we include additional losses that enforce temporal and spatial consistency of long rollouts. These include: (i) a temporal Sobolev loss matching finite-difference time derivatives, (ii) a spatial gradient loss enforcing edge and structure consistency, and (iii) a spectral loss computed in the Fourier domain to penalize frequency and phase mismatches. All physics terms are aggregated into a single auxiliary loss $\mathcal{L}_{\mathrm{phys}}$. More details can be found in Appendix C.
Overall objective.
The final training objective is given by
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{recon}}+\alpha\,\mathcal{L}_{\mathrm{pred}}+\beta\,\mathcal{L}_{\mathrm{latent}}+\lambda_{\mathrm{phys}}\,\mathcal{L}_{\mathrm{phys}},$ (14)
where $\alpha$ and $\beta$ control the relative contributions of the rollout prediction and latent consistency losses, respectively, and $\lambda_{\mathrm{phys}}$ weights the optional physics-informed regularization terms. More details about how the hyperparameters were chosen can be found in C.
[…]
Figure 1: Architecture overview. The history encoder and present encoder (top) feed into the latent state $z_{t_{0}}$, followed by a dynamics rollout and decoding at each step.
4.3 Inference and Integration via Matrix Exponentiation
A major advantage of restricting the latent dynamics to a linear Koopman form is that, at inference time, we can leverage the analytical solution of the learned system through matrix exponentiation. This bypasses iterative ODE solvers entirely and yields efficient long-horizon forecasting. Given the linear system in (8), the exact solution for any future state $z_{\tau}$ given $z_{0}$ is defined by the matrix exponential $z_{\tau}=\exp(\mathbf{K}_{\text{cont}}\tau)z_{0}$. This formulation allows us to predict the state at an arbitrary future time $\tau$ in a single computational step, independent of the training step size $\Delta t$. The matrix exponential $\exp(\mathbf{M})$ is defined by the power series:
$\exp(\mathbf{M})=\sum_{n=0}^{\infty}\frac{\mathbf{M}^{n}}{n!}=\mathbf{I}+\mathbf{M}+\frac{\mathbf{M}^{2}}{2!}+\frac{\mathbf{M}^{3}}{3!}+\dots$ (15)
5 Datasets
We compare our model with the benchmarking Autoregressive Conditional Diffusion Models (ACDM) (Kohl et al., 2026) and use the same dataset they generate for training and evaluation as well. We use two scenarios of flow in the dataset, an incompressible wake flow and a transonic cylinder flow with shock waves.
Incompressible Wake Flow.
This dataset consists of a fully developed incompressible Karman vortex street behind a cylindrical obstacle with varying Reynolds numbers $Re\in[100,1000]$, with 91 simulations in total. For training, both ACDM and our model utilize simulations within the range $Re\in[200,900]$. There are two distinct test sets: $Inc_{low}$, covering $Re\in[100,180]$, and $Inc_{high}$, covering $Re\in[920,1000]$, with a prediction horizon of $T=60$ steps. While the raw datasets employ a time step of $\Delta t=0.05\text{s}$, we follow the baseline protocol and subsample the trajectories by a factor of 2, resulting in an effective time step of $\Delta t=0.1\text{s}$ for model training and evaluation.
Transonic Cylinder Flow.
Transonic flows are more chaotic, exhibiting the formation of shock waves that interact with the flow. The dataset consists of 40 sequences of fully developed compressible Karman vortex streets at a fixed Reynolds number $Re=10,000$, with Mach numbers varying across the range $Ma\in[0.50,0.90]$. Both models are trained on sequences with $Ma\in[0.53,0.63]\ \cup\ [0.69,0.90]$. Two of the testsets $Tra_{ext}$ and $Tra_{int}$ represent simulations with $Ma\in[0.50,0.52]$ and $Ma\in[0.66,0.68]$ respectively with rollout steps $T=60$, and another long rollout case of about 8 vortex shedding periods, $Tra_{long}$, has $Ma\in[0.64,0.65]$ and $T=240$. Similar to the incompressible case, the raw simulation interval is $\Delta t=0.05\text{s}$, which is subsampled to $\Delta t=0.1\text{s}$ for our experiments.
6 Experiments
In this section, we present a comprehensive evaluation of our proposed method. First, we detail the Experimental Setup, contrasting our training strategies with ACDM baseline to highlight key methodological differences. Subsequently, we report the Results, assessing the model’s performance on incompressible and transonic flows.
6.1 Experimental Setup
While we utilize the same underlying data as the ACDM baseline, we introduce specific methodological refinements to the training protocol to maximize data efficiency and ensure a rigorous comparison. All models were trained to forecast the next $N=8$ states conditioned on the present one and the previous one.
Data Loading Strategy:
Standard training of the ACDM baseline utilizes non-overlapping sequences to form training batches. This approach discards a significant portion of valid temporal transitions. In contrast, similar to established data augmentation practices in autoregressive surrogate modeling (Brandstetter et al., 2022; Wu et al., 2022), we adopt an exhaustive sliding-window strategy. By sliding the temporal window forward by a single time step, we generate heavily overlapping batches (i.e., if a sequence begins at $t_{i}$, the next valid training sequence begins at $t_{i+1}$). The sliding-window strategy increases the number of observed temporal transitions, improving supervision of the latent dynamics. Importantly, this does not introduce additional simulation data, but reuses existing trajectories to provide denser temporal coverage. As a result, the model learns a more consistent latent representation while maintaining a fair comparison with baselines.
Loss Function Formulation:
While diffusion models like ACDM can be modified to include physics-informed guidance during the sampling phase, their standard training objective inherently relies on a probabilistic MSE derived from the Gaussian diffusion assumption. In contrast, our deterministic framework naturally accommodates a direct, composite loss function (as defined in (14)) during the primary training loop. This allows us to strictly enforce physics-informed constraints—such as Sobolev norms and spectral alignment—leading to sharper structural predictions without requiring auxiliary guidance models at inference.
Training Regimes:
ACDM baseline for the Incompressible Flow dataset was trained on the fully developed regime. While effective for capturing stationary statistics, this protocol excludes the initial transient phase where the flow develops from rest. To ensure a rigorous comparison of generalization capabilities, we define two distinct evaluation settings:
  • Stable-Only Training: Models are trained only on the fully developed regime. This includes the original baseline (ACDM), its no-noise-conditional alternative (ACDMncn), and our standard model (Continuous KAE).
  • Full-Dataset Training: Models are trained on the complete evolution history, including the transient initialization. We retrained the baseline from scratch on the full dataset (Full retrained ACDM) and compare it against our Full Continuous KAE model.
Note that for the Transonic Flow, all models (baseline and ours) were trained on the full dataset by default.
6.2 Results
We structure our evaluation into five main parts: a quantitative analysis comparing our method against the baselines, a validation of the learned integrator against the analytical matrix exponential, an assessment of temporal consistency across varying resolutions, and a final study on computational efficiency.
We compare against ACDM as a strong generative baseline, alongside autoregressive and Koopman-based models, to evaluate trade-offs between expressivity, stability, and efficiency. Benchmarking against it allows us to explicitly characterize the spectrum of trade-offs in PDE forecasting: the high-frequency synthesis and short-term accuracy of diffusion models versus the temporal flexibility, extreme long-horizon stability, and computational efficiency of continuous-time Koopman theory. Generative models achieve lower short-term error by modeling fine-scale stochastic structure. In contrast, our method produces smoother predictions due to the linear structure of the latent dynamics, which can lead to slightly higher error at short horizons.
We also compare training and inference efficiency across models. While all methods are trained on the same underlying simulation data, their computational costs differ significantly due to their respective optimization objectives and architectures. Diffusion-based models require iterative noise conditioning and sampling, resulting in longer training and inference times. In contrast, our deterministic formulation enables substantially faster convergence and evaluation.
In our experiments, the continuous-time KAE converges in approximately 200 epochs, compared to 1000 epochs for the diffusion baseline, corresponding to wall-clock training times of approximately $24$ hours versus $72$ hours on a single NVIDIA RTX 4090 GPU. We further observe a significant reduction in inference cost, as our model requires only a single forward pass per timestep.
Quantitative Analysis and Phase Alignment.
Tables 1 and 3 detail the quantitative performance across regimes, revealing a clear trade-off between short-term accuracy and long-horizon stability. On the Incompressible Flow dataset (Table 1), the Full Continuous KAE effectively captures the smoother dynamics, outperforming the Full retrained ACDM in both extrapolation tasks ($Inc_{low}$ and $Inc_{high}$). The deterministic nature of our model handles the predictable vortex shedding highly effectively.
However, in the highly chaotic Transonic Flow regimes (Table 3), the differing strengths of the frameworks become apparent. For the interpolation task ($Tra_{int}$) and near-extrapolation ($Tra_{ext}$), in highly chaotic transonic regimes, diffusion models achieve lower short-term MSE and LSiM by synthesizing fine-scale stochastic features. The diffusion baseline excels here because it can stochastically generate the sharp, high-frequency shock waves characteristic of compressible flows. However, this advantage diminishes over long horizons, where our deterministic, $L_{2}$-optimized KAE exhibits a spectral bias, tending to smooth these chaotic, high-frequency features (frequency-space metrics quantifying this smoothing effect are provided in Appendix E). To better visualize this, difference maps between the ground truth and predictions are included in Appendix F, showing that KAE errors are localized entirely around the shock fronts, whereas ACDM errors are more diffusely distributed.
This smoothing becomes advantageous during the long-horizon rollout ($Tra_{long}$, $T=240$). Over extended autoregressive steps, the stochastic nature of ACDM leads to error accumulation and structural drift, resulting in an MSE of $22.2\times 10^{-3}$. In contrast, the linear latent dynamics remain stable over long horizons. This stability arises from the spectral structure of the learned Koopman operator. As visualized in Appendix E.3, the eigenvalue distribution of the learned operator across several conditioning parameters lies predominantly in the left half of the complex plane, corresponding to dissipative latent dynamics. This structure naturally suppresses unstable growth modes during rollout and helps explain the improved long-horizon stability observed in our experiments. By restricting the latent dynamics to a mathematically stable evolution, our KAE degrades gracefully and achieves a significantly lower MSE of $17.0\times 10^{-3}$, demonstrating superior robustness for extreme long-term forecasting.
For probabilistic models, direct comparison to a single reference trajectory is insufficient, as trajectories naturally decorrelate over time even for accurate simulations (hu2020). To further quantify the temporal degradation of the models, Figure 2 presents the spatial Pearson correlation coefficient between the predicted flow fields and the reference simulations over the 240-step $Tra_{long}$ horizon. While both models maintain high correlation initially, ACDM gradually loses phase alignment with the reference vortex shedding, resulting in a steep drop in correlation and high variance across samples. The Continuous KAE demonstrates a notably slower rate of correlation decay and a tighter variance band, confirming that the linear constraints of the Koopman latent space effectively preserve the underlying temporal phase. A deeper investigation into the spectral properties of this smoothing effect, alongside an extreme 1000-step stress test, is provided in Appendix E.
Figure 2: Spatial Pearson correlation over a 240-step rollout in the Transonic regime ($Tra_{long}$). The Continuous KAE maintains significantly more stable structural alignment compared to the unconstrained diffusion baseline.
Integrator Analysis.
We compare trajectories generated using numerical RK4 integration with the analytical matrix exponential solution derived in Section 4.3. While RK4 performs iterative rollouts with a fixed timestep $\Delta t=0.1\,\mathrm{s}$, the analytical formulation allows direct computation of the latent state as $z(\tau)=\exp(K\tau)z_{0}$, for any target time $\tau$. Figure 4 and Table 3 show close agreement between the two approaches, confirming that the learned dynamics are well approximated by the linear generator. This property enables highly efficient long-horizon forecasting, as future states can be computed in a single step without sequential rollout.
Temporal Consistency and Generalization.
A key advantage of our continuous-time KAE is its ability to generalize to arbitrary temporal resolutions without retraining. Despite being trained on a fixed timestep $\Delta t=0.1s$, the model supports arbitrary integration step sizes during inference. We validate this capability in Figure 5, which displays rollouts generated using three distinct integration step sizes, $\Delta t\in\{0.05s,0.1s,0.2s\}$. For each case, the number of rollout steps is adjusted to align with the same set of physical time instants (up to $t=2.0s$). The first row also shows the results produced using the analytical exponential form. The perfect alignment confirms that the learned dynamics are robust to discretization changes.
(a) Vorticity for $Inc_{high}$ with $Re=1000$
(b) Pressure for $Tra_{ext}$ with $Ma=0.50$
Figure 3: Qualitative validation rollouts for extrapolation regimes. (Left) Vorticity prediction for the incompressible wake flow in the high-Reynolds-number regime ($Inc_{high},Re=1000$). (Right) Pressure prediction for the transonic cylinder flow in the low-Mach extrapolation regime ($Tra_{ext},Ma=0.50$).
Method $Inc_{low}$ $Inc_{high}$
MSE $(\times 10^{-4})$ LSiM $(\times 10^{-2})$ MSE $(\times 10^{-5})$ LSiM $(\times 10^{-2})$
$ACDM_{ncn}$ $175.1\pm 14.3$ $73.3\pm 4.4$ $2348.3\pm 291.3$ $82.5\pm 4.0$
ACDM $8.5\pm 19.2$ $12.6\pm 14.9$ $223.3\pm 614.2$ $10.4\pm 21.7$
Full retrained ACDM $1.4\pm 1.7$ $6.2\pm 4.6$ $2.4\pm 1.9$ $1.9\pm 0.5$
Continuous KAE $9.2\pm 20.9$ $13.2\pm 16.3$ $213.7\pm 457.6$ $12.6\pm 22.3$
Full Continuous KAE $\mathbf{1.0\pm 1.3}$ $\mathbf{5.2\pm 4.4}$ $\mathbf{1.4\pm 2.6}$ $\mathbf{1.0\pm 0.7}$
Table 1: Quantitative comparison on incompressible wake flow extrapolation regimes ($Inc_{low}$ and $Inc_{high}$). Performance is reported using MSE and LSiM (where lower values indicate better performance), averaged over rollout timesteps. Simple ACDM and ACDMncn, as well as Continuous KAE, were trained only on the fully developed regime, whereas the Full retrained models were trained on the full dataset.
Architecture Avg. Step (ms) Mean VRAM (MB)
ResNet-m2 $3.67\pm 0.04$ $188.0$
Dil-ResNet-m2 $3.46\pm 0.02$ $\mathbf{178.6}$
FNO-16 $1.17\pm 0.01$ $184.1$
FNO-32 $1.17\pm 0.00$ $183.9$
UNet-m2 $6.19\pm 0.09$ $183.7$
UNet-m8 $6.16\pm 0.01$ $184.1$
TF-Enc $0.60\pm 0.25$ $3448.6$
TF-MGN $0.69\pm 0.01$ $3498.0$
TF-VAE $0.30\pm 0.01$ $13749.9$
Refiner-R4 $10.31\pm 0.02$ $642.4$
KAE $\mathbf{1.04\times 10^{-3}\pm 10^{-4}}$ $2751.3$
ACDM $41.77\pm 0.01$ $659.2$
$\text{ACDM}_{ncn}$ $41.70\pm 0.06$ $649.2$
Table 2: Comprehensive runtime comparison between ACDM and continuous KAE across a 240-step rollout. The table is split into two parts for clarity: the top table presents the breakdown of KAE runtimes by model components (Encode, Decode) alongside the total 240-step rollout time, allowing a direct comparison of per-component computational overhead. The bottom table reports derived per-step metrics, including the average computation time per step and the total average per-step time across the rollout, which are more directly indicative of the model’s efficiency in long-horizon predictions. All values are reported as mean $\pm$ standard deviation in milliseconds (ms) across multiple independent runs.
Method $Inc_{low}$ $Inc_{high}$ $Tra_{ext}$ $Tra_{int}$ $Tra_{long}*$
MSE $(\times 10^{-4})$ LSiM $(\times 10^{-2})$ MSE $(\times 10^{-5})$ LSiM $(\times 10^{-2})$ MSE $(\times 10^{-3})$ LSiM $(\times 10^{-1})$ MSE $(\times 10^{-3})$ LSiM $(\times 10^{-1})$ MSE $(\times 10^{-3})$ LSiM $(\times 10^{-1})$
ResNet $10\pm 9.1$ $17\pm 7.8$ $16\pm 3.0$ $5.9\pm 1.6$ $2.3\pm 0.9$ $1.4\pm 0.2$ $1.8\pm 1.0$ $\mathbf{1.0\pm 0.3}$ $24.2\pm 4.6$ $7.6\pm 1.7$
ResNet-dil $1.6\pm 1.8$ $7.7\pm 5.5$ $2.6\pm 0.7$ $1.7\pm 1.0$ $1.2\pm 0.3$ $1.7\pm 1.4$ $\mathbf{1.0\pm 0.5}$ $5.7\pm 2.1$ $22.0\pm 2.4$ $5.5\pm 2.3$
$\text{FNO}_{16}$ $2.8\pm 3.1$ $8.8\pm 7.1$ $8.9\pm 3.8$ $2.5\pm 1.2$ $4.8\pm 1.2$ $3.4\pm 1.1$ $5.5\pm 2.6$ $2.6\pm 1.1$ $20.8\pm 2.0$ $11.5\pm 1.1$
$\text{FNO}_{32}$ $160\pm 50$ $80\pm 5.4$ $1k\pm 140$ $57\pm 4.9$ $4.9\pm 1.9$ $3.6\pm 0.9$ $6.8\pm 3.4$ $3.1\pm 1.1$ Diverged Diverged
$\text{TF}_{MGN}$ $5.7\pm 4.3$ $13\pm 6.4$ $10\pm 2.9$ $3.5\pm 0.4$ $3.9\pm 1.0$ $1.8\pm 0.3$ $6.3\pm 4.4$ $2.2\pm 0.7$ $18.9\pm 4.5$ $4.0\pm 0.2$
$\text{TF}_{Enc}$ $1.5\pm 1.7$ $6.3\pm 4.2$ $8.7\pm 3.8$ $0.6\pm 0.3$ $\mathbf{1.0\pm 0.3}$ $3.3\pm 1.2$ $1.8\pm 0.3$ $6.2\pm 4.2$ $22.2\pm 3.9$ $3.8\pm 0.4$
$\text{TF}_{VAE}$ $5.4\pm 5.5$ $13\pm 7.2$ $4.1\pm 1.4$ $4.1\pm 0.9$ $2.4\pm 0.2$ $7.2\pm 3.0$ $2.7\pm 0.6$ $11\pm 5.1$ $20.6\pm 2.1$ $4.0\pm 0.2$
U-Net $1.0\pm 1.1$ $5.8\pm 3.2$ $2.7\pm 0.6$ $2.6\pm 0.6$ $3.1\pm 2.1$ $3.9\pm 2.8$ $2.3\pm 2.0$ $3.3\pm 2.8$ $30.3\pm 6.1$ $9.1\pm 1.2$
[…]
$\text{U-Net}_{tn}$ $1.0\pm 1.0$ $5.6\pm 3.1$ $0.9\pm 0.6$ $1.5\pm 0.6$ $4\pm 0.8 1.1$$1\pm 0.3$ $1.8\pm 1.1$ $1.0\pm 0.4$ $22.4\pm 3.9$ $3.9\pm 0.4$
Refiner $1.3\pm 1.4$ $7.1\pm 4.2$ $3.5\pm 2.2$ $2.5\pm 1.0$ $5.4\pm 2.1$ $2.3\pm 0.5$ $7.1\pm 2.1$ $3.0\pm 1.7$ Diverged $8.8\pm 3.3$
$\text{ACDM}_{ncn}$ $0.9\pm 0.8$ $6.6\pm 2.7$ $5.7\pm 2.7$ $3.6\pm 1.2$ $4.1\pm 1.9$ $1.9\pm 0.6$ $2.8\pm 1.3$ $1.7\pm 0.4$ $22.8\pm 3.8$ $8.2\pm 3.0$
ACDM $1.7\pm 2.2$ $6.9\pm 5.6$ $0.8\pm 0.4$ $1.0\pm 0.3$ $2.3\pm 1.4$ $1.3\pm 0.3$ $2.7\pm 2.1$ $1.3\pm 0.6$ $22.6\pm 4.0$ $\mathbf{3.8\pm 0.4}$
Continuous KAE $1.3\pm 1.7$ $6.1\pm 4.8$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
Exponential Continuous KAE $1.3\pm 1.7$ $6.1\pm 4.8$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
Table 3: Quantitative comparison across incompressible and transonic flow regimes, including extrapolation ($Inc_{low},Inc_{high},Tra_{ext},Tra_{int}$) and long-horizon rollout ($Tra_{long}$). Metrics reported are MSE and LSiM. Results for $Tra_{long}$ are computed by the authors, as this setting is not reported in the original ACDM paper.
Efficiency.
[…]
(a) Vorticity for $Inc_{high}$ with $Re=1000$
(b) Pressure for $Tra_{ext}$ with $Ma=0.50$
Figure 4: Comparison between numerical RK4 integration and the analytical matrix exponential solution of the learned continuous-time latent dynamics. Results are shown for (Left) incompressible flow vorticity at $Re=1000$ and (Right) transonic flow pressure at $Ma=0.50$.
Figure 5: Direct comparison of inference results using the matrix exponentiation and integrator at different time steps. All columns display the flow’s evolution at the same physical time and each row represents a different evolution scheme: first one is the direct matrix exponentiation, while the subsequent ones use the RK4 integrator with different time steps.
Asymptotic Stability and Dissipation.
To test the absolute limits of the learned operator, we subjected the models to an extreme 1000-step rollout ($T=1000$), shown in Appendix E.2. Over such extended horizons, autoregressive data-space models typically suffer from catastrophic numerical divergence, generating non-physical artifacts. In contrast, our Continuous KAE degrades gracefully. Because the learned latent dynamics are fundamentally governed by a linear Koopman operator, the system is asymptotically stable; the eigenvalues of the operator capture the natural dissipative properties of the fluid. As the rollout progresses towards $t=1000$, the high-frequency transient modes naturally decay, and the prediction smoothly diffuses toward the steady-state base flow. While this results in a loss of fine-scale turbulent texture (quantified by the spectral analysis in Appendix E), it leads to bounded predictions over long horizons, making it highly reliable for extreme long-term structural forecasting. This behavior is consistent with the non-autoregressive nature of our model: because the latent evolution is computed directly via matrix exponentiation, errors do not compound across timesteps.
Figure 6: Visual snapshots of the flow field over an extreme 1000-step rollout in the Transonic regime. While the Continuous KAE smoothly diffuses the flow into a stable limit cycle without blowing up, the unconstrained autoregressive diffusion baseline (ACDM) eventually compounds stochastic errors until the physical structure collapses into numerical noise.
7 Discussion
In this work, we presented a continuous-time Koopman Autoencoder that effectively decouples the learning of latent dynamics from the temporal resolution of the training data. Our results on incompressible and transonic flows demonstrate that this approach provides an alternative to probabilistic methods like ACDM. They reveal a trade-off between expressivity and dynamical consistency. While generative models capture high-frequency stochastic features, they rely on autoregressive sampling that can lead to instability over long horizons. In contrast, structured latent dynamics impose a strong inductive bias that favors stable and efficient evolution, at the cost of reduced expressivity. This suggests that hybrid approaches combining structured dynamics with expressive residual modeling may offer a promising direction for future work.
Accuracy vs. Efficiency Trade-off. While diffusion models such as ACDM excel at capturing high-frequency texture through stochastic refinement, they incur a significant computational cost during inference due to iterative denoising steps. Our method, by contrast, utilizes a deterministic ODE solver in the latent space. As shown in Table 1, our continuous formulation achieves competitive MSE and LSiM scores, often within the error margins of ACDM, while being two orders of magnitude faster during inference. This efficiency is beneficial for applications requiring fast evaluation.
Convergence Speed. As a result of the exhaustive data loading and the deterministic nature of our model, we achieve convergence significantly faster than the baseline. Our models are trained for 200 epochs, whereas ACDM requires 1000 epochs.
Temporal Generalization. A key advantage of the continuous-time formulation is robustness to temporal resolution. Our model maintains physical consistency when queried at time intervals different from training. Crucially, when the learned latent dynamics are evolved using the analytical matrix exponentiation, predictions remain stable and closely match those obtained via numerical integration of RK4. This demonstrates that the model learns a valid continuous-time generator of the dynamics, enabling zero-shot evaluation at arbitrary temporal resolutions. In contrast, discrete-time baselines typically fail under such temporal mismatches or require retraining, indicating reliance on step-to-step mappings rather than an underlying dynamical operator.
Spectral Bias and Extreme Long-Horizon Stability. As reflected in our transonic results, the deterministic nature of our continuous-time KAE introduces a spectral bias: it tends to smooth fine-scale, chaotic spatial features. Our frequency analysis (detailed in Appendix E) confirms this, showing a faster energy drop-off at high spatial wavenumbers compared to the diffusion-based ACDM, which synthesizes textural details more accurately. However, the temporal frequency analysis reveals the KAE’s critical advantage: it captures the dominant vortex shedding frequencies with minimal variance. Rather than a strictly negative artifact, this physical low-pass filtering effect is exactly what guarantees the model’s stability. When subjected to an extreme $T=1000$ step rollout, ACDM’s unconstrained generation compounds high-frequency stochastic errors until the physical structure collapses. In contrast, our Continuous KAE remains strictly bounded by its linear latent dynamics, gracefully decaying high-frequency transients to settle into a stable, physically accurate limit cycle without numerical divergence. To address the spectral smoothing while retaining this extreme stability, future hybrid approaches could utilize the continuous-time Koopman state as a mathematically rigorous base flow. A lightweight, conditional diffusion or generative adversarial decoder could then be applied exclusively at the final inference step to stochastically recover fine-scale turbulent textures without risking autoregressive error accumulation.
Limitations. Despite these successes, we observe that the spectral bias of the $L_{2}$ reconstruction loss can lead to the smoothing of fine-scale turbulent features over very long horizons, a common issue in deterministic autoencoders. Additionally, while the Koopman operator provides interpretability, the assumption of global linearity (even with residual corrections) remains a constraint for highly non-linear phenomena like shock wave propagation in the transonic regime.
8 Conclusion
We introduced a continuous-time Koopman Autoencoder for fluid dynamics forecasting, modelling latent evolution as an ODE governed by a parameter-conditioned linear generator. This formulation combines the interpretability of Koopman theory with the flexibility of neural differential equations, enabling accurate, zero-shot temporal-resolution-invariant predictions. Extensive benchmarks on incompressible and transonic flows confirm that our approach offers massive computational efficiency and extreme long-horizon stability compared to highly expressive autoregressive diffusion models.
Future work will scale this approach to three-dimensional turbulence—made computationally viable by our low-rank parameter adaptation—and incorporate explicit conservation laws into the latent ODE to further improve physical fidelity. Ultimately, this work demonstrates that enforcing structured, continuous-time linear dynamics in latent spaces provides a scalable, efficient, and robust alternative for complex PDE forecasting where long-term stability is critical.
This work demonstrates that enforcing simple linear structure in latent space can yield robust and scalable models for dynamical systems. By combining Koopman theory with continuous-time dynamics, we obtain models that are temporally consistent, computationally efficient, and stable over long horizons. These results suggest that structured latent representations provide a viable alternative to highly expressive autoregressive models, particularly in regimes where stability and generalization are critical. More broadly, this work highlights the value of combining physical structure with learned representations to build scalable and reliable models for complex dynamical systems.
Impact Statement
This paper introduces a surrogate model that significantly accelerates fluid dynamics simulations. By reducing the reliance on computationally expensive numerical solvers (DNS/LES), this method facilitates more energy-efficient workflows in engineering and climate science. This contributes to lowering the carbon footprint associated with high-performance computing, with no foreseeable negative ethical or societal risks.
Software and Data
Datasets are publicly available at https://github.com/tum-pbs/autoreg-pde-diffusion/tree/main at the training and evaluation resolution of $128\times 64$ ($\sim$146 GB).
Appendix A Data Details
All models were trained on the public data from (Kohl et al., 2026). For a thorough comparison, we used the exact same split in terms of train/evaluation. The incompressible model was trained on all the sequences with Reynolds numbers in the $[200,900]$ interval with a step of $10$, and evaluated on the $Inc_{low}=[100,200]$, $Inc_{high}=[900,1000]$ extreme intervals. The transonic model was also trained on the sequences with Mach numbers in the $0.53,0.54,...,0.62,0.63]\cup[0.69,0.70,...,0.89,0.90]$ and evaluated on the sequences with $Tra_{ext}=[0.50,0.51,0.52]$, $Tra_{int}=[0.66,0.67,0.68]$ for $60$ time steps and $Tra_{long}=[0.64,0.65]$ for $240$ time steps. The main difference in training and loading the data, as mentioned in Section 6.1 is we treat them in an exhaustive manner. We follow the same stride (a sub-sampling factor of $2$), but we load the data such that the previous sequence’s prediction can be initial conditions for the following sequence.
Moreover, we used a different structure. The original Diffusion Model was using a simple tensor with multiple channels, each of them representing a different variable ($v_{x}$, $v_{y}$, etc.). For expressiveness we used TensorDicts, which are PyTorch’s (Paszke et al., 2019) version of dictionaries, where the values are tensors. Each of the channels from the original formulation was translated into a specific variable.
Appendix B Implementation and Model Details
The entire model is implemented and trained using PyTorch (Paszke et al., 2019). All wights in the architecture were optimized using AdamW (Loshchilov and Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate follows a Cosine Warm-up schedule, linear increasing for the first $20$ epochs from $0.0$ to $5\times 10^{-4}$, then decreasing on a cosine curve to $10^{-5}$ for the following $180$ epochs, making it $200$ epochs in total for training, with a batch size of $64$.
We also follow an end-to-end training procedure, learning the encoder, operator and decoder simultaneously. While testing different architectures and training schemes, we have experimented with curriculum learning too. As the model can be rather complex, especially for large latent dimensionality, while the data, more specifically the transonic flow, can be very chaotic and rough, we tried training the Auto-Encoder’s core (encoder and decoder) for the first $50$ epochs alone, freezing the operator’s side, as we first wanted to build a smooth manifold for the data. For the following $100$ epochs, we have frozen the encoder and decoder and only trained the operator. Finally, for the remaining epochs, we fine-tuned them together. However, this proved challenging as regularizing the latent space was not as straightforward as expected, leaving the linear operator to struggle.
Our model follows a modular encoderlatentdynamicsdecoderencoder–latent–dynamics–decoder design, combining convolutional feature extraction with physics-conditioned Koopman operators for latent-space temporal evolution. The architecture is specifically designed for stability, interpretability, and robustness in strongly conditioned fluid dynamics regimes.
B.1 Encoder Architecture
The spatial encoder is a convolutional residual encoder operating on input fields augmented with explicit coordinate information. For each input snapshot of shape $C\times H\times W$, two additional coordinate channels corresponding to normalized spatial coordinates are concatenated, resulting in $C+2$ input channels. This coordinate injection enables the model to reason about absolute spatial position without relying on implicit convolutional biases.
The encoder backbone begins with an initial $3\times 3$ convolution, followed by a sequence of pre-activation residual blocks (He et al., 2016) with Group Normalization (Wu and He, 2018) and SiLU activations. Each resolution level consists of one downsampling residual block with stride $2$, and one refinement residual block with stride $1$.
We use three resolution levels with channel widths $[64,128,256]$, leading to an effective downsampling factor of $2^{3}=8$. All convolutional and linear layers in the encoder employ spectral normalization, which we found crucial for stabilizing gradients in downstream Koopman dynamics learning.
At the lowest spatial resolution, we apply a Convolutional Block Attention Module (CBAM) (Woo et al., 2018), consisting of sequential channel-wise and spatial attention. This allows the encoder to focus on dynamically active regions such as vortices while suppressing background noise.
The resulting feature map is flattened and projected into a latent representation via a linear layer. This latent representation significantly improves numerical stability of the Koopman operator. When late fusion conditioning is enabled, the flattened spatial features are concatenated with a physics embedding before the final projection.
Conditioning and Physics Parameter Embeddings Physical parameters (e.g. Reynolds number, Mach number, forcing terms) are incorporated through radial basis function (RBF)–style expansions, mapping low-dimensional scalars into higher-dimensional embeddings. The expansion type is configurable and shared across encoder, decoder, and dynamics modules. During training, small Gaussian noise can be injected into the conditioning variables to improve robustness to discretization artifacts and sparse parameter sampling.
B.1.1 Temporal History Encoder
For scenarios involving multiple input timesteps, we introduce a history encoder that aggregates temporal context before latent evolution. Each timestep is independently encoded using the convolutional encoder backbone, producing a sequence of latent vectors. These vectors are then processed by a Transformer encoder (Vaswani et al., 2017) with sinusoidal positional encodings. The Transformer output is averaged across time to obtain a single latent context vector, which serves as the initial condition for the Koopman dynamics. This design allows the model to capture temporal correlations without explicitly unrolling convolutional operations over time.
B.2 Decoder Architecture
The decoder mirrors the encoder structure in reverse, reconstructing spatial fields from the latent state. The latent vector is first linearly expanded and reshaped into a low-resolution feature map. Upsampling proceeds through a sequence of resolution levels $[256,128,64]$, each consisting of a conditioned residual block, and a resize–convolution upsampling block (nearest-neighbor upsampling followed by a $3\times 3$ convolution).
To integrate physical parameters, we employ Adaptive Group Normalization (AdaGN) within the decoder residual blocks. Here, scale and shift parameters of the normalization layers are modulated by the physics embeddings, enabling strong, spatially uniform conditioning while preserving convolutional locality. Dropout is applied inside conditioned residual blocks for regularization, using element-wise dropout to avoid suppressing entire physical channels. The final output is produced via a $3\times 3$ convolution mapping back to the original number of physical variables.
B.3 Koopman Dynamics Module
Latent evolution is governed by a physics-conditioned Koopman operator implemented in continuous-time form. We consider two Koopman parameterizations:
  • Linear mode, where the dynamics are governed by a base linear operator with low-rank, condition-dependent updates predicted by a hyper-network (LoRA-style adaptation) (Hu et al., 2022).
  • MLP-augmented mode, where a linear operator is supplemented by a nonlinear residual MLP to capture weakly nonlinear effects.
Condition dependence is handled via a hyper-network that predicts operator modifications from the physics embeddings. Stability is explicitly enforced through dissipative initialization, spectral constraints, and, in continuous-time settings, stable numerical integration.
For continuous dynamics, we model $\frac{d\mathbf{z}}{dt}=K({\phi})\mathbf{z}$, and integrate using a fourth-order Runge–Kutta (RK4) scheme by default, with optional implicit midpoint or Radau IIA solvers for stiff regimes. All implicit solvers internally upcast to $float32$ to ensure numerical robustness.
Appendix C Training Objective and Loss Function Details
To ensure that the learned latent dynamics are not only accurate in terms of Euclidean error but also physically consistent, stable, and topologically faithful to the fluid flows, we employ a composite loss function. The total objective $\mathcal{L}_{\text{total}}$ is a weighted sum of reconstruction accuracy, temporal rollout consistency, latent space regularization, and physics-informed constraints:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{recon}}+\alpha\mathcal{L}_{\text{pred}}+\beta\mathcal{L}_{\text{latent}}+\lambda_{\text{phys}}\mathcal{L}_{\text{phys}}$ (16)
C.1 Loss Weighting Strategy
The scalar weights $\alpha,\beta$, and $\lambda_{phys}$ are critical for balancing the trade-offs between spatial reconstruction fidelity, temporal stability, and adherence to theoretical Koopman constraints.
  • Reconstruction Base ($\mathcal{L}_{recon}$): Kept at a unit weight, this anchors the overall scale of the loss. It is heavily prioritized in the early epochs to ensure the encoder and decoder construct a valid spatial manifold before temporal dynamics are strictly enforced.
  • Rollout Prediction ($\alpha$): This weight dictates the model’s resistance to autoregressive error accumulation. A sufficiently high $\alpha$ forces the Koopman operator to learn a globally stable trajectory rather than a greedy one-step mapping. However, if $\alpha$ dominates too early in training, the network struggles to converge on the underlying spatial representations.
  • Latent Consistency ($\beta$): This parameter controls the strictness of the theoretical Koopman constraints (detailed below). Balancing $\beta$ is crucial: if it is too low, the latent space evolves non-linearly, risking divergence during long rollouts; if it is too high, it over-constrains the autoencoder, resulting in overly smoothed spatial reconstructions that fail to capture complex flow features.
  • Physics Regularization ($\lambda_{phys}$): Because spatial gradients and spectral amplitudes operate on different numerical scales than standard pixel-wise MSE, $\lambda_{phys}$ scales these high-frequency penalties. It acts as a structural fine-tuning mechanism to sharpen edges and correct phase shifts once the base dynamics are established.
Below, we detail the mathematical formulation of the latent and physics-informed components implemented in our framework.
C.2 Latent Space Consistency
To ensure the latent manifold respects the theoretical properties of the Koopman operator, we enforce structural constraints via $\mathcal{L}_{\text{latent}}$. This includes three specific sub-terms:
1. Forward-Backward Linearity Consistency. A valid Koopman operator must be invertible. If $\mathbf{K}(\Delta t)$ propagates the state forward, $\mathbf{K}(-\Delta t)$ must recover the previous state. This is crucial for dissipative systems to prevent the operator from learning a trivial ”shrink-to-zero” solution.
$\mathcal{L}_{\text{lin}}=\underbrace{\|\mathbf{K}(z_{t},\Delta t)-z_{t+1}\|_{2}^{2}}_{\text{Forward}}+\underbrace{\|\mathbf{K}(z_{t+1},-\Delta t)-z_{t}\|_{2}^{2}}_{\text{Backward}}.$ (17)
2. Directional Cosine Similarity. To decouple magnitude errors (decay) from directional errors (dynamics), we enforce alignment between the predicted latent vector update and the true encoder trajectory:
$\mathcal{L}_{\text{cos}}=1-\frac{z_{t+1}\cdot\mathcal{E}(x_{t+1})}{\|z_{t+1}\|\|\mathcal{E}(x_{t+1})\|}.$ (18)
3. Energy Conservation Regularization. Ideally, the ”energy” (norm) of the latent state should evolve smoothly. We penalize abrupt changes in the latent norm to encourage smooth trajectories:
$\mathcal{L}_{\text{energy}}=\left(\|z_{t+1}\|_{2}-\|z_{t}\|_{2}\right)^{2}.$ (19)
The total latent loss is given by $\mathcal{L}_{\text{latent}}=\mathcal{L}_{\text{lin}}+w_{cos}\mathcal{L}_{\text{cos}}+\mathcal{L}_{\text{energy}}$.
C.3 Physics-Informed Loss
Standard MSE losses often result in ”pacing” errors (phase shift) or blurred edges. We mitigate this using a physics-informed loss $\mathcal{L}_{\text{phys}}$ comprising Sobolev norms and spectral analysis.
Temporal Sobolev Loss (Velocity Matching). Enforces consistency in the time-derivative (velocity) of the flow:
$\mathcal{L}_{\text{time}}=\left\|\frac{\partial\hat{x}}{\partial t}-\frac{\partial x}{\partial t}\right\|_{2}^{2}\approx\left\|(\hat{x}_{t+1}-\hat{x}_{t})-(x_{t+1}-x_{t})\right\|_{2}^{2}.$ (20)
Spatial Sobolev Loss (Structure Matching). Enforces consistency in spatial gradients to preserve sharp edges (e.g., shock waves):
$\mathcal{L}_{\text{space}}=\|\nabla_{x}\hat{x}-\nabla_{x}x\|_{2}^{2}+\|\nabla_{y}\hat{x}-\nabla_{y}x\|_{2}^{2}.$ (21)
Spectral Consistency Loss. To correct phase errors and ensure the model captures the correct shedding frequencies, we compute the loss in the frequency domain using the Fast Fourier Transform ($\mathcal{F}$). This term penalizes discrepancies in both amplitude (energy spectrum) and phase:
$\mathcal{L}_{\text{spectral}}=\||\mathcal{F}(\hat{x})|-|\mathcal{F}(x)|\|_{1}+\|\text{Re}(\mathcal{F}(\hat{x}))-\text{Re}(\mathcal{F}(x))\|_{2}^{2}+\|\text{Im}(\mathcal{F}(\hat{x}))-\text{Im}(\mathcal{F}(x))\|_{2}^{2}.$ (22)
Appendix D Error Analysis
In this appendix, we provide a detailed quantitative and qualitative comparison between our method and ACDM, focusing on accuracy, stability, and robustness across flow regimes. Beyond aggregate error metrics, we emphasize distributional and temporal error characteristics to better understand when and why our approach outperforms diffusion-based forecasting.
We visualize prediction errors using a combination of time-resolved error curves, error bar plots, and violin plots. Error bars report the mean and standard deviation of normalized $\ell_{2}$ errors across test trajectories, highlighting both average performance and variability. Violin plots are used to capture the full error distribution over space and time, revealing differences in tail behavior and robustness that are not visible from mean metrics alone.
Across all evaluated scenarios, our model exhibits consistently lower variance in prediction error, indicating improved stability under long-horizon rollouts. In particular, the error distributions produced by ACDM show heavier tails, corresponding to occasional but severe prediction failures, whereas our Koopman-based model yields tighter, more concentrated distributions. This effect is especially pronounced in transonic regimes, where diffusion-based models tend to suffer from mode collapse or over-smoothing.
We further include timestep-sensitivity plots, where the same trained model is evaluated under varying rollout step sizes. These visualizations demonstrate that our continuous-time latent dynamics maintain accuracy across a wide range of timesteps, while ACDM performance degrades significantly when evaluated outside its training resolution. This highlights a key advantage of the continuous-time formulation in terms of temporal generalization.
Finally, spatial error maps are presented to illustrate qualitative differences in failure modes. Our model localizes errors primarily in dynamically active regions such as shocks or vortex cores, while maintaining low background error. In contrast, ACDM exhibits spatially diffuse errors that accumulate over time, consistent with the absence of an explicit latent evolution operator.
(a) Low Reynolds regime
(b) High Reynolds regime
Figure 7: Field-averaged mean squared error (MSE) comparison between KAE and ACDM for incompressible flows. Error bars denote one standard deviation across test trajectories. Our method consistently achieves lower error and reduced variance, with the performance gap widening in the high-Re regime where chaotic dynamics dominate.
(a) Low Reynolds regime
(b) High Reynolds regime
Figure 8: Temporal evolution of field-wise MSE for incompressible flows. Errors are shown on a logarithmic scale. KAE maintains stable error growth over long horizons, while ACDM exhibits accelerated error accumulation, indicative of compounding stochastic prediction drift.
(a) Low Reynolds regime
(b) High Reynolds regime
Figure 9: Field-wise MSE as a function of Reynolds number for incompressible flows. Velocity components and pressure are shown separately. KAE exhibits consistently smoother error scaling with Reynolds number, while ACDM shows increased sensitivity and variance, particularly in pressure-dominated regimes.
(a) Low Reynolds number
(b) High Reynolds number
Figure 10: Error distributions under low and high Reynolds number regimes. While both models perform comparably at low Reynolds numbers, ACDM exhibits pronounced heavy-tailed error distributions at high Reynolds numbers. In contrast, KAE maintains controlled variance, demonstrating superior robustness in turbulent regimes.
Appendix E Extended Stability and Spectral Analysis
To thoroughly address the accuracy-stability trade-offs between deterministic latent dynamics and stochastic diffusion, this section provides an extended analysis of the models’ spectral biases and their behavior under extreme autoregressive rollouts.
E.1 Spectral Bias and Frequency Smoothing
As established in turbulence research (dryden1943), even highly accurate numerical solvers eventually decorrelate from a target simulation over extended timeframes due to the chaotic nature of fluid dynamics. Therefore, we evaluate the temporal and spatial frequency spectra to measure whether the predicted trajectories statistically match the physical characteristics of the reference simulation. Figure 11 presents the temporal frequency of the vertical motion ($v_{y}$) and the spatial wavenumber of the horizontal motion ($v_{x}$) evaluated downstream.
The temporal analysis (Figure 11, left) demonstrates that the Continuous KAE accurately identifies and locks onto the primary vortex shedding frequency (the dominant energy peak), exhibiting a very tight variance band. ACDM also captures this primary frequency, though its stochastic formulation introduces slightly more variance.
In the spatial domain (Figure 11, right), ACDM accurately synthesizes the high-frequency turbulent textures. Conversely, the Continuous KAE exhibits a steeper energy drop-off at high wavenumbers. This confirms that our model suppresses high-frequency chaotic structures in favor of a stable limit cycle. It acts as a low-pass physical filter, discarding the chaotic, unpredictable high-frequency turbulent cascade to guarantee that the primary, macro-scale shedding frequencies are preserved with near-zero variance.
Figure 11: Temporal and spatial frequency analysis on a sequence from $Tra_{long}$. The shaded area represents the 5th to 95th percentile. The Continuous KAE successfully captures the dominant physical frequencies but attenuates high-frequency turbulent noise compared to the stochastic baseline.
E.2 Extreme Long-Horizon Stability ($T=1000$)
To test the absolute limits of the learned latent dynamics and investigate potential numerical divergence, we subjected both models to an extreme 1000-step autoregressive rollout (as visualized in Figure 6).
As shown in Figure 12, the danger of ACDM’s textural hallucinations becomes catastrophic over extreme horizons. The stochastic nature of the diffusion model leads to severe phase divergence: its relative $L_{2}$ error spikes erratically with massive variance, and its spatial correlation completely collapses as the high-frequency details compound into unphysical numerical noise.
In contrast, the Continuous KAE degrades gracefully. Because the latent space is governed by a globally stable linear Koopman operator, the system is asymptotically bounded. The transient, high-frequency modes naturally decay, leaving only the stable base flow. While fine-scale textures diffuse out over time, the model maintains a bounded $L_{2}$ error and a highly stable periodic correlation with the reference simulation, proving its reliability for robust long-term forecasting without structural collapse.
(a) Spatial Correlation (1000 steps)
(b) Relative $L_{2}$ Error (1000 steps)
Figure 12: Quantitative metrics over an extreme 1000-step rollout. ACDM exhibits severe instability and variance, while the Continuous KAE remains strictly bounded by its linear latent dynamics.
E.3 Latent Dynamics Eigenvalue Spectrum
To better understand the exceptional long-horizon stability of the continuous-time Koopman Autoencoder, we analyze the spectral properties of the learned latent dynamics. For a linear continuous-time system governed by $\frac{dz}{dt}=\mathbf{K}z$, the system is strictly stable if the real parts of all eigenvalues of $\mathbf{K}$ are negative.
As shown in Figure 13, we plot the eigenvalue distribution of the learned generator matrix $\mathbf{K}_{\text{cont}}(\phi)$ across several conditioning parameters. The spectrum lies predominantly in the left half of the complex plane ($Re(\lambda)<0$). This mathematically corresponds to dissipative latent dynamics, which naturally suppress high-frequency unstable growth modes. Consequently, any numerical errors introduced during latent projection or temporal rollout naturally decay over time rather than compounding exponentially, ensuring bounded and stable predictions even over extreme time horizons.
Figure 13: Eigenvalue spectrum of the learned Koopman generator across several conditioning parameters. The majority of eigenvalues lie in the stable region ($Re(\lambda)<0$), indicating dissipative latent dynamics.
Appendix F Spatial Error Distribution and Difference Maps
To analyze the physical nature of the prediction errors, we provide absolute spatial difference maps across all evaluated regimes. While raw MSE metrics often favor stochastic models in chaotic flows, these visual distributions highlight the trade-off between textural detail and structural coherence.
F.1 Comparative Analysis: Continuous KAE vs. Diffusion Models
The fundamental difference between the Continuous KAE and the ACDM diffusion model lies in how they handle high-frequency information and temporal evolution:
  • Continuous KAE (Deterministic Stability): The KAE is constrained by a linear Koopman operator in a continuous latent space. This structure ensures asymptotic stability, meaning errors do not grow exponentially over time. However, the deterministic $L_{2}$ objective leads to a ”spectral bias” where the model prefers to smooth out high-frequency spatial discontinuities to ensure global structural alignment.
  • Diffusion Models (Stochastic Detail): ACDM uses a stochastic process to synthesize fine-scale turbulent textures. While this allows the model to match the sharp ’look’ of the ground truth, it lacks the rigid mathematical constraints of the KAE. Consequently, minor stochastic errors compound over time, leading to widespread spatial noise and phase divergence.
F.2 Visual Error Localization across All Regimes
The difference maps reveal two distinct archetypes of model failure. Errors generated by the Continuous KAE are highly localized along sharp discontinuities such as shock fronts and vortex edges. In contrast, errors from the ACDM baseline are more diffusely distributed throughout the spatial domain.
(a) Transonic Interpolation
(b) Transonic Extrapolation
Figure 14: Spatial error distribution in Transonic regimes. KAE errors are concentrated precisely at the sharp shock fronts, while ACDM exhibits broader spatial noise.
Figure 15: Spatial error distribution in the long-rollout regime. The KAE maintains structural stability with localized errors, while ACDM shows diffuse error growth across the wake.
(a) Incompressible (Low Reynolds)
(b) Incompressible (High Reynolds)
Figure 16: Comparison of error magnitudes across incompressible flow regimes. Note the higher error localization on the obstacle boundary for the KAE.